# The Certainty Ratio $\mathcal{C_{\rho}}$: a novel metric for assessing the reliability of classifier predictions

### v1

@@ Abstract
Evaluating the performance of classifiers is critical in machine learning, particularly in high–stakes applications where the reliability of predictions can significantly impact decision–making. Traditional performance measures, such as accuracy and F–score, often fail to account for the uncertainty inherent in classifier predictions, leading to potentially misleading assessments. This paper introduces the Certainty Ratio ($C_{\rho}$), a novel metric designed to quantify the contribution of confident (certain) versus uncertain predictions to any classification performance measure. By integrating the Probabilistic Confusion Matrix ($CM^{\star}$) and decomposing predictions into certainty and uncertainty components, $C_{\rho}$ provides a more comprehensive evaluation of classifier reliability. Experimental results across 26 datasets and multiple classifiers, including Decision Trees, Naïve–Bayes, 3–Nearest Neighbors, and Random Forests, demonstrate that $C_{\rho}$ reveals critical insights that conventional metrics often overlook. These findings emphasize the importance of incorporating probabilistic information into classifier evaluation, offering a robust tool for researchers and practitioners seeking to improve model trustworthiness in complex environments.
@@ 5 Uncertainty Matrix
$\begin{bmatrix}2.3&0.2&0.5\\ 0.5&1.1&0.4\\ 0&0.9&0.1\\ \end{bmatrix}=\begin{bmatrix}2.3&0&0\\ 0.5&0.8&0\\ 0&0.9&0\\ \end{bmatrix}+\begin{bmatrix}0&0.2&0.5\\ 0.1&0.3&0.4\\ 0&0&0.1\\ \end{bmatrix}$
@@
The matrix $V$ collects the contribution to the $CM^{\star}$ of potentially correct predictions made by the classifier $\rho$. It represents the certainty of the classifier, as it assigns a class to a test example with higher probability than to other classes. From the matrix $V$ any probabilistic measure can be calculated. For example, the $Acc^{\star}_{v}=\frac{2.3+0.8+0}{6}=0.517$.
@@
The matrix $U$ collects the contribution to the probabilistic confusion matrix $CM^{\star}$ from potentially incorrect predictions made by the classifier $\rho$. It represents the uncertainty of the classifier, as the probabilities present in this matrix correspond to classes different from the predicted class of $\rho$. From the matrix $U$, various probabilistic measure can be calculated. For example, the probabilistic accuracy is given by $Acc^{\star}_{u}=\frac{0+0.3+0.1}{6}=0.07$.
@@ 6 Probabilistic measures
$Acc=\frac{1}{n}\sum_{i=1}^{k}\alpha_{i,i}$
where $n$ is the number of test examples (obtained as the sum of all the elements in the matrix), $k$ is the number of classes (number of rows or columns of the matrix), and $\forall i\in\{1,\dots,k\}$, $\alpha_{i,i}\in CM_{k,k}$.
@@
From the $CM^{\star}$ matrix, the accuracy $Acc^{\star}=\frac{3.5}{6}=0.583$. It is evident that $Acc^{\star}$ can be decoomposed as $Acc^{\star}=Acc_{u}^{\star}+Acc_{v}^{\star}=0.517+0.07$.
@@
$\phi(CM^{\star}_{k,k})=\phi_{v}(V_{k,k})+\phi_{u}(U_{k,k})$ (14)
@@
$Acc^{\star}=\frac{1}{n}\sum_{i=1}^{k}\beta_{i,i}=\frac{1}{n}\sum_{i=1}^{k}v_{i,i}+\frac{1}{n}\sum_{i=1}^{k}u_{i,i}$ (15)
where $\forall i\in\{1,\dots,k\}$, $\beta_{i,i}\in CM^{\star}_{k,k}$, $v_{i,i}\in V$ and $u_{i,i}\in U$.
@@
$\mathcal{C_{\rho}}=\frac{\phi_{v}(V_{k,k})}{\phi_{v}(V_{k,k})+\phi_{u}(U_{k,k})}=\frac{\phi_{v}(V_{k,k})}{\phi(CM^{\star}_{k,k})}$ (17)
@@
Given the values of accuracy $Acc^{\star}_{v}=0.517$ and $Acc^{\star}_{u}=0.07$, the certainty ratio $\mathcal{C_{\rho}}$ is:
$\mathcal{C_{\rho}}=\frac{0.517}{0.517+0.07}=0.88$
This means that 88% of predictions are attributed to certainty, while 12% are due to uncertainty.
@@ 7 Experimental analysis
australian credit 0.664 0.632 0.557 0.074 5.8 88.2 0.599
cardiotocography 0.718 0.675 0.625 0.051 3.0 92.5 0.595
chronic kidney 0.769 0.736 0.682 0.054 4.8 92.6 0.712
customer churn 0.865 0.841 0.811 0.031 4.1 96.4 0.686
data banknote 0.999 1.000 0.999 0.000 0.0 100.0 1.000
ecoli 0.854 0.808 0.786 0.022 3.3 97.3 0.585
fertility diagnosis 0.890 0.817 0.810 0.007 10.9 99.2 0.503
hepatitis c 0.265 0.246 0.143 0.103 12.5 58.2 0.238
landsat satellite 0.912 0.896 0.876 0.019 1.3 97.8 0.872
leaf 0.585 0.546 0.485 0.061 3.3 88.9 0.518
magic gamma 0.800 0.769 0.727 0.042 4.7 94.5 0.716
movement libras 0.789 0.753 0.711 0.041 2.1 94.5 0.736
musk v1 0.866 0.826 0.786 0.041 3.9 95.1 0.813
musk v2 0.969 0.946 0.938 0.008 3.4 99.2 0.905
parkinson 0.836 0.824 0.779 0.044 4.1 94.6 0.748
phishing websites 0.942 0.932 0.916 0.016 1.6 98.3 0.921
planning relax 0.648 0.582 0.502 0.080 14.6 86.2 0.460
rice cammeo 0.890 0.858 0.837 0.021 3.5 97.5 0.841
room occupancy 0.997 0.997 0.996 0.001 0.1 99.9 0.987
sonar 0.812 0.793 0.755 0.038 2.7 95.2 0.774
speaker accent 0.818 0.763 0.721 0.042 5.3 94.6 0.705
vertebral column 0.806 0.799 0.747 0.052 1.6 93.5 0.729
vowel 0.983 0.961 0.955 0.006 1.0 99.4 0.954
winequality red 0.533 0.492 0.400 0.092 6.7 81.3 0.249
winequality white 0.501 0.469 0.371 0.099 7.4 79.0 0.252
yeast 0.561 0.502 0.432 0.070 5.9 86.1 0.489
Mean 0.780 0.748 0.706 0.043 4.5 92.3 0.676
Table 3: Results for Naïve–Bayes.
@@
australian credit 0.823 0.818 0.794 0.024 0.5 97.1 0.786
cardiotocography 0.445 0.440 0.408 0.032 0.6 92.6 0.568
chronic kidney 0.952 0.951 0.948 0.004 0.1 99.6 0.957
customer churn 0.731 0.723 0.701 0.022 1.2 97.0 0.788
data banknote 0.839 0.810 0.764 0.046 2.9 94.4 0.743
ecoli 0.792 0.775 0.749 0.026 1.7 96.6 0.521
fertility diagnosis 0.850 0.793 0.769 0.024 9.8 97.0 0.468
hepatitis c 0.256 0.250 0.082 0.168 3.8 32.8 0.291
landsat satellite 0.797 0.796 0.791 0.005 0.1 99.4 0.782
leaf 0.724 0.719 0.702 0.017 1.0 97.6 0.709
magic gamma 0.728 0.721 0.695 0.026 2.9 96.4 0.628
movement libras 0.644 0.641 0.632 0.009 0.7 98.6 0.634
musk v1 0.655 0.657 0.648 0.009 0.2 98.6 0.673
musk v2 0.671 0.671 0.666 0.005 0.0 99.3 0.716
parkinson 0.703 0.703 0.696 0.007 1.0 99.0 0.771
phishing websites 0.627 0.642 0.615 0.027 2.1 95.8 0.682
planning relax 0.648 0.568 0.472 0.095 23.3 83.2 0.448
rice cammeo 0.920 0.917 0.909 0.008 0.3 99.1 0.904
room occupancy 0.950 0.937 0.930 0.007 1.1 99.3 0.797
sonar 0.668 0.674 0.652 0.022 0.7 96.8 0.675
speaker accent 0.574 0.542 0.497 0.045 2.7 91.7 0.558
vertebral column 0.816 0.772 0.727 0.045 3.9 94.2 0.671
vowel 0.685 0.566 0.494 0.072 4.7 87.3 0.516
winequality red 0.565 0.483 0.385 0.098 5.8 79.7 0.298
winequality white 0.447 0.400 0.288 0.112 5.9 72.1 0.253
yeast 0.197 0.205 0.168 0.037 3.1 81.8 0.408
Mean 0.681 0.661 0.622 0.038 3.1 91.4 0.625
@@
australian credit 0.823 0.817 0.800 0.017 0.7 97.9 0.777
cardiotocography 0.867 0.848 0.841 0.008 1.0 99.1 0.747
chronic kidney 0.970 0.965 0.965 0.000 0.6 100.0 0.942
customer churn 0.940 0.936 0.929 0.007 0.5 99.2 0.855
data banknote 0.983 0.980 0.980 0.000 0.3 100.0 0.962
ecoli 0.830 0.798 0.792 0.006 2.5 99.2 0.484
fertility diagnosis 0.810 0.784 0.764 0.020 4.2 97.4 0.511
hepatitis c 0.240 0.246 0.219 0.026 1.2 89.3 0.241
landsat satellite 0.865 0.858 0.852 0.006 0.4 99.3 0.809
leaf 0.615 0.599 0.582 0.016 1.6 97.3 0.570
magic gamma 0.824 0.816 0.806 0.010 0.9 98.8 0.768
movement libras 0.653 0.636 0.617 0.020 1.4 96.9 0.605
musk v1 0.775 0.772 0.762 0.010 0.4 98.8 0.743
musk v2 0.968 0.966 0.964 0.002 0.2 99.8 0.925
parkinson 0.862 0.847 0.840 0.007 2.2 99.2 0.736
phishing websites 0.959 0.952 0.945 0.006 0.9 99.3 0.931
planning relax 0.626 0.619 0.602 0.016 1.2 97.4 0.537
rice cammeo 0.892 0.888 0.880 0.008 0.6 99.1 0.855
room occupancy 0.996 0.995 0.995 0.000 0.1 100.0 0.971
sonar 0.731 0.726 0.716 0.009 0.9 98.7 0.697
speaker accent 0.681 0.661 0.648 0.013 1.7 98.1 0.549
vertebral column 0.803 0.786 0.778 0.008 1.6 98.9 0.686
vowel 0.793 0.782 0.770 0.012 0.6 98.5 0.756
winequality red 0.607 0.586 0.566 0.020 2.0 96.6 0.320
winequality white 0.593 0.572 0.551 0.021 1.9 96.4 0.319
yeast 0.516 0.495 0.464 0.031 1.5 93.7 0.357
Mean 0.778 0.767 0.755 0.012 1.2 98.0 0.679
@@
australian credit 0.868 0.789 0.749 0.039 8.0 95.0 0.699
cardiotocography 0.898 0.773 0.750 0.023 5.5 97.0 0.631
chronic kidney 0.992 0.960 0.958 0.003 3.4 99.7 0.914
customer churn 0.957 0.927 0.914 0.013 3.0 98.6 0.805
data banknote 0.992 0.979 0.976 0.003 1.4 99.7 0.948
ecoli 0.866 0.750 0.719 0.031 6.8 95.9 0.453
fertility diagnosis 0.880 0.807 0.786 0.021 10.9 97.4 0.516
hepatitis c 0.250 0.248 0.081 0.167 3.0 32.7 0.290
landsat satellite 0.920 0.838 0.817 0.021 4.5 97.5 0.758
leaf 0.756 0.532 0.485 0.047 5.1 91.2 0.482
magic gamma 0.881 0.806 0.772 0.034 8.6 95.8 0.712
movement libras 0.814 0.571 0.532 0.039 7.1 93.2 0.515
musk v1 0.943 0.799 0.777 0.022 14.6 97.2 0.696
musk v2 0.980 0.947 0.941 0.006 3.9 99.4 0.845
parkinson 0.908 0.833 0.804 0.029 8.8 96.5 0.701
phishing websites 0.968 0.939 0.930 0.009 3.0 99.1 0.895
planning relax 0.698 0.586 0.505 0.081 25.0 86.1 0.462
rice cammeo 0.925 0.889 0.872 0.017 3.5 98.1 0.845
room occupancy 0.998 0.993 0.992 0.001 0.3 99.9 0.947
sonar 0.827 0.677 0.603 0.074 15.7 89.1 0.591
speaker accent 0.812 0.594 0.551 0.043 12.7 92.7 0.471
vertebral column 0.839 0.773 0.722 0.051 5.2 93.4 0.647
vowel 0.974 0.706 0.699 0.007 9.1 99.0 0.613
winequality red 0.699 0.569 0.493 0.075 10.3 86.7 0.321
winequality white 0.697 0.558 0.482 0.076 11.8 86.4 0.321
yeast 0.639 0.485 0.413 0.072 8.5 85.2 0.388
Mean 0.845 0.743 0.705 0.039 7.7 92.4 0.633
@@
The experimental analysis section is designed to evaluate the effectiveness of the probabilistic confusion matrix across various classifiers and datasets. The study involves 26 datasets from the UCI Machine Learning Repository [11], varying in the number of samples (100 to 19,020), number of variables (4 to 166), and number of classes (2 to 30) (see Table 1). The performance of 3–Nearest Neighbors (3–NN), Naïve Bayes (NB), Decision Trees (DT), and Random Forests (RF) classifiers is assessed using both traditional and probabilistic performance measures. Two classification performance measures were selected: accuracy and the IMCP curve [12, 13]. The choice of the IMCP curve is motivated by the fact that this measure is calculated from the probabilities returned by the classifier. Unlike $Acc$ or other popular measures such as F–score or Matthews correlation coefficient, the area under the IMCP curve implicitly considers the probabilistic nature of predictions.
@@
Tables 2, 3, 4 and 5 show, for each dataset (row), several columns: Dataset name, accuracy ($Acc$), probabilistic accuracy ($Acc^{\star}$), certainty accuracy ($Acc_{v}^{\star}$), uncertainty accuracy ($Acc_{u}^{\star}$), divergence ($div$) in percentage, certainty ratio ($\mathcal{C}_{\rho}$) in percentage, and the area under the IMCP curve (IMCP).
@@
The 3–Nearest Neighbor classifier exhibits stable behavior, with low divergence of about 4.5%, resulting in a small difference between $Acc=0.780$ and $Acc^{\star}=0.748$, and a certainty ratio $\mathcal{C}_{3-NN}=92.3\%$. These results indicate that 3–NN maintains consistent performance with relatively low uncertainty.
@@
The Naïve–Bayes classifier shows similar behavior to the 3–NN classifier but with generally lower accuracy and uncertainty. The mean accuracy is $Acc=0.681$, and the certainty ratio is 91.4%. This suggests that Naïve–Bayes often provides confident predictions, but these predictions are not always as accurate.
@@
The Random Forest classifier achieves the highest accuracy ($Acc=0.845$), suggesting it is the best classifier. However, it also has a relatively high divergence of 7.7% and a certainty ratio of 92.4%, indicating that a significant portion of its performance comes from uncertain predictions. This implies that despite high accuracy, RF may not always be the most reliable.
@@
In contrast, the Decision Tree classifier exhibits a very high certainty ratio of 98%, indicating robust and reliable performance. Although its accuracy ($Acc=0.778$) is slightly lower than RF, its lower divergence and higher certainty ratio make it a more dependable classifier. The Decision Tree’s results for the IMCP are similar to or even better than those of Naïve–Bayes and Random Forest, underscoring its reliability.
@@
The probabilistic accuracy ($Acc_{v}^{\star}$) tends to be lower than traditional accuracy ($Acc$), highlighting the overestimation issue when uncertainty is ignored. The IMCP curve, which considers probabilistic predictions, consistently shows values closer to $Acc_{v}^{\star}$ than to $Acc$, supporting the argument that probabilistic measures offer a more realistic assessment of classifier performance.
@@
In summary, relying solely on accuracy is not advisable when the certainty ratio is not high (very close to 100%). More broadly, we should avoid relying on any classification performance measure when the certainty ratio indicates a significant contribution of uncertainty in the confusion matrix. Therefore, it is more appropriate to analyze the probabilistic confusion matrix to assess the reliability of a performance measure, guided by the certainty ratio $\mathcal{C}_{\rho}$. This approach ensures that the evaluation of a classifier accounts not only for overall performance but also for the level of confidence and certainty in its predictions.
@@
3–NN 0.780 0.748 0.706 0.043 4.5 92.3 0.676
Decision Tree 0.778 0.767 0.755 0.012 1.2 98.0 0.679
Naive Bayes 0.681 0.661 0.622 0.038 3.1 91.4 0.625
Random Forest 0.845 0.743 0.705 0.039 7.7 92.4 0.633

### v2

@@ Abstract
Evaluating the performance of classifiers is critical in machine learning, particularly in high–stakes applications where the reliability of predictions can significantly impact decision–making. Traditional performance measures, such as accuracy and F–score, often fail to account for the uncertainty inherent in classifier predictions, leading to potentially misleading assessments. This paper introduces the Certainty Ratio ($C_{\rho}$), a novel metric designed to quantify the contribution of confident (certain) versus uncertain predictions to any classification performance measure. By integrating the Probabilistic Confusion Matrix ($CM^{\star}$) and decomposing predictions into certainty and uncertainty components, $C_{\rho}$ provides a more comprehensive evaluation of classifier reliability. Experimental results across 21 datasets and multiple classifiers, including Decision Trees, Naïve–Bayes, 3–Nearest Neighbors, and Random Forests, demonstrate that $C_{\rho}$ reveals critical insights that conventional metrics often overlook. These findings emphasize the importance of incorporating probabilistic information into classifier evaluation, offering a robust tool for researchers and practitioners seeking to improve model trustworthiness in complex environments.
@@ 5 Uncertainty Matrix
$\begin{bmatrix}2.3&0.2&0.5\\ 0.5&1.1&0.4\\ 0&0.9&0.1\\ \end{bmatrix}=\begin{bmatrix}2.3&0&0\\ 0.4&0.8&0\\ 0&0.9&0\\ \end{bmatrix}+\begin{bmatrix}0&0.2&0.5\\ 0.1&0.3&0.4\\ 0&0&0.1\\ \end{bmatrix}$
@@
The matrix $V$ collects the contribution to the $CM^{\star}$ of potentially correct predictions made by the classifier $\rho$. It represents the certainty of the classifier, as it assigns a class to a test example with higher probability than to other classes. From the matrix $V$ any probabilistic measure can be calculated.
@@
The matrix $U$ collects the contribution to the probabilistic confusion matrix $CM^{\star}$ from potentially incorrect predictions made by the classifier $\rho$. It represents the uncertainty of the classifier, as the probabilities present in this matrix correspond to classes different from the predicted class of $\rho$. From the matrix $U$, various probabilistic measure can be calculated.
@@ 6 Probabilistic measures
$Acc=\frac{\sum_{i=1}^{k}\alpha_{i,i}}{\sum_{i,j=1,1}^{k,k}\alpha_{i,j}}$
where $k$ is the number of classes (number of rows or columns of the matrix), and $\forall i,j\in\{1,\dots,k\}$, $\alpha_{i,j}\in CM_{k,k}$.
@@
From the $CM^{\star}$ matrix, the accuracy $Acc^{\star}=\frac{3.5}{6}=0.583$.
@@
$\phi(CM^{\star}_{k,k})=\tau\left(\phi_{v}(V_{k,k}),\phi_{u}(U_{k,k})\right)$ (14)
where $\tau$ is a function.
@@
$Acc^{\star}=\lambda_{v}Acc^{\star}_{v}+\lambda_{u}Acc^{\star}_{u}$ (15)
where
$\lambda_{v}=\frac{\sum_{i,j=1,1}^{k,k}v_{i,j}}{\sum_{i,j=1,1}^{k,k}(u_{i,j}+v_{i,j})}$ (16)
$\lambda_{u}=\frac{\sum_{i,j=1,1}^{k,k}u_{i,j}}{\sum_{i,j=1,1}^{k,k}(u_{i,j}+v_{i,j})}$ (17)
$Acc^{\star}_{v}=\frac{\sum_{i=1}^{k}v_{i,i}}{\sum_{i,j=1,1}^{k,k}v_{i,j}}$ (18)
$Acc^{\star}_{u}=\frac{\sum_{i=1}^{k}u_{i,i}}{\sum_{i,j=1,1}^{k,k}u_{i,j}}$ (19)
where $v_{i,j}\in V_{k,k}$, and $u_{i,j}\in U_{k,k}$.
In Example 10, $Acc^{\star}=0.583$, which can be decomposed as follows:
$Acc^{\star}=\frac{4.4}{6}\cdot\frac{3.1}{4.4}+\frac{1.6}{6}\cdot\frac{0.4}{1.6}=0.73\cdot 0.705+0.27\cdot 0.25=0.583$
The coefficients $\lambda_{v}$ and $\lambda_{u}$ are extremely important. $\lambda_{v}$ measures the ratio of probabilities used in the certainty matrix $V$, and $\lambda_{u}$ measures the ratio of probabilities used in the uncertainty matrix $U$. In the example, $\lambda_{v}=0.73$, which means that about 73% of predictions were made with confidence (this does not mean that they were accurate). In contrast, the system assigned about 27% of probabilities to predictions without confidence (and some might even be accurate by chance).
@@
$\mathcal{C_{\rho}}=\frac{\phi_{v}(V_{k,k})}{\phi_{v}(V_{k,k})+\phi_{u}(U_{k,k})}$ (21)
@@
Considering that $\phi_{v}(V_{k,k})=Acc^{\star}_{v}$ and $\phi_{u}(U_{k,k})=Acc^{\star}_{u}$:
$Acc^{\star}_{v}=\frac{2.3+0.8}{4.4}=0.705\quad Acc^{\star}_{u}=\frac{0.3+0.1}{1.6}=0.25$ (22)
the certainty ratio $\mathcal{C_{\rho}}$ is:
$\mathcal{C_{\rho}}=\frac{0.705}{0.705+0.25}=0.738$
This means that when measuring accuracy for classifier $\rho$ about 74% of predictions are attributed to certainty, while 26% are due to uncertainty.
@@ 7 Experimental analysis
australian credit 0.672 0.636 0.689 0.401 6.0 63.2 0.603
cardiotocography 0.719 0.680 0.768 0.297 3.8 72.2 0.601
customer churn 0.866 0.843 0.886 0.360 3.6 71.3 0.687
data banknote 1.000 1.000 1.000 0.000 0.0 100.0 1.000
fertility diagnosis 0.880 0.813 0.880 0.117 10.4 90.7 0.547
hepatitis c 0.256 0.244 0.259 0.225 11.6 53.5 0.236
landsat satellite 0.911 0.894 0.930 0.327 1.3 74.0 0.868
magic gamma 0.801 0.770 0.822 0.363 4.6 69.4 0.716
movement libras 0.808 0.767 0.848 0.261 5.5 77.4 0.754
musk v1 0.861 0.828 0.890 0.362 4.6 71.4 0.817
musk v2 0.965 0.943 0.973 0.217 3.3 81.8 0.902
parkinson 0.852 0.832 0.873 0.395 5.2 70.3 0.747
phishing websites 0.949 0.940 0.960 0.367 1.0 72.4 0.933
planning relax 0.622 0.579 0.630 0.411 14.7 61.8 0.472
rice cammeo 0.888 0.856 0.902 0.290 3.5 75.7 0.839
room occupancy 0.997 0.996 0.997 0.362 0.1 76.0 0.987
sonar 0.821 0.803 0.843 0.398 4.1 69.6 0.785
speaker accent 0.833 0.769 0.878 0.225 7.2 80.8 0.702
vertebral column 0.816 0.802 0.853 0.409 5.4 68.7 0.726
vowel 0.974 0.952 0.982 0.246 2.0 81.2 0.944
winequality red 0.508 0.497 0.553 0.346 4.6 61.5 0.252
Mean 0.809 0.783 0.829 0.304 4.9 73.5 0.720
Table 3: Results for Naïve-Bayes.
@@
australian credit 0.787 0.790 0.802 0.524 2.5 61.9 0.764
cardiotocography 0.592 0.589 0.607 0.373 1.4 62.0 0.656
customer churn 0.732 0.731 0.740 0.493 0.4 60.1 0.799
data banknote 0.843 0.810 0.871 0.361 3.7 70.9 0.745
fertility diagnosis 0.850 0.794 0.861 0.236 10.4 81.9 0.506
hepatitis c 0.240 0.248 0.241 0.251 11.2 48.6 0.290
landsat satellite 0.798 0.795 0.801 0.347 0.3 70.3 0.782
magic gamma 0.727 0.721 0.737 0.450 2.9 62.1 0.631
movement libras 0.633 0.631 0.640 0.364 2.1 66.6 0.630
musk v1 0.735 0.735 0.739 0.470 1.4 65.5 0.737
musk v2 0.839 0.838 0.841 0.437 0.2 65.9 0.795
parkinson 0.713 0.709 0.715 0.525 1.4 60.5 0.771
phishing websites 0.604 0.616 0.613 0.726 1.7 46.0 0.662
planning relax 0.660 0.571 0.659 0.336 23.4 66.8 0.452
rice cammeo 0.911 0.906 0.917 0.379 0.5 71.0 0.893
room occupancy 0.964 0.956 0.969 0.281 0.9 77.7 0.853
sonar 0.669 0.674 0.678 0.556 2.7 56.6 0.673
speaker accent 0.574 0.546 0.627 0.225 6.1 73.8 0.561
vertebral column 0.823 0.790 0.854 0.341 4.7 72.0 0.688
vowel 0.657 0.549 0.708 0.233 5.8 75.3 0.503
winequality red 0.546 0.473 0.565 0.293 6.3 65.9 0.293
Mean 0.709 0.689 0.723 0.390 4.3 65.8 0.652
@@
australian credit 0.851 0.851 0.851 0.000 0.0 100.0 0.848
cardiotocography 0.855 0.855 0.856 0.227 0.2 86.2 0.798
customer churn 0.940 0.940 0.943 0.508 0.4 66.9 0.878
data banknote 0.986 0.986 0.986 0.000 0.0 100.0 0.986
fertility diagnosis 0.760 0.755 0.759 0.000 0.7 100.0 0.613
hepatitis c 0.258 0.258 0.258 0.000 0.0 100.0 0.257
landsat satellite 0.861 0.861 0.861 0.000 0.0 100.0 0.839
magic gamma 0.818 0.818 0.818 0.000 0.0 100.0 0.800
movement libras 0.706 0.706 0.706 0.000 0.0 100.0 0.715
musk v1 0.771 0.771 0.771 0.000 0.0 100.0 0.77
musk v2 0.971 0.971 0.971 0.000 0.0 100.0 0.944
parkinson 0.872 0.872 0.872 0.000 0.0 100.0 0.822
phishing websites 0.963 0.963 0.968 0.479 0.3 67.1 0.961
planning relax 0.665 0.665 0.665 0.000 0.0 100.0 0.586
rice cammeo 0.876 0.876 0.876 0.000 0.0 100.0 0.873
room occupancy 0.995 0.995 0.995 0.000 0.0 100.0 0.985
sonar 0.725 0.725 0.725 0.000 0.0 100.0 0.725
peaker accent 0.690 0.690 0.690 0.000 0.0 100.0 0.599
vertebral column 0.781 0.781 0.781 0.000 0.0 100.0 0.701
vowel 0.819 0.819 0.819 0.000 0.0 100.0 0.819
winequality red 0.623 0.623 0.623 0.000 0.0 100.0 0.368
Mean 0.799 0.799 0.800 0.058 0.0 96.2 0.757
@@
australian credit 0.872 0.789 0.893 0.238 8.6 79.0 0.699
cardiotocography 0.895 0.769 0.924 0.124 5.8 88.2 0.625
customer churn 0.956 0.925 0.968 0.241 3.2 80.3 0.800
data banknote 0.993 0.978 0.995 0.123 1.7 90.0 0.944
fertility diagnosis 0.860 0.802 0.876 0.236 11.0 81.1 0.546
hepatitis c 0.258 0.252 0.260 0.248 7.5 50.9 0.292
landsat satellite 0.916 0.837 0.942 0.170 4.4 84.7 0.758
magic gamma 0.882 0.805 0.902 0.235 8.9 79.3 0.711
movement libras 0.836 0.570 0.894 0.077 10.3 92.1 0.515
musk v1 0.899 0.751 0.917 0.165 15.8 84.9 0.652
musk v2 0.978 0.943 0.984 0.136 4.0 87.9 0.839
parkinson 0.913 0.835 0.931 0.201 9.2 83.0 0.702
phishing websites 0.972 0.950 0.980 0.214 2.2 82.2 0.913
planning relax 0.698 0.578 0.695 0.293 27.5 70.5 0.460
rice cammeo 0.922 0.889 0.936 0.267 3.3 77.9 0.846
room occupancy 0.998 0.993 0.998 0.079 0.4 93.1 0.946
sonar 0.818 0.695 0.853 0.274 14.3 76.0 0.605
speaker accent 0.805 0.589 0.855 0.127 14.3 87.1 0.465
vertebral column 0.842 0.77 0.876 0.269 7.2 76.7 0.648
vowel 0.973 0.696 0.983 0.028 9.7 97.3 0.605
winequality red 0.698 0.572 0.745 0.223 10.3 76.9 0.320
Mean 0.856 0.761 0.877 0.189 8.6 81.9 0.661
@@
The experimental analysis section is designed to evaluate the effectiveness of the probabilistic confusion matrix across various classifiers and datasets. The study involves 21 datasets from the UCI Machine Learning Repository [11], varying in the number of samples (100 to 19,020), number of variables (4 to 166), and number of classes (2 to 15) (see Table 1). The performance of 3–Nearest Neighbors (3–NN), Naïve Bayes (NB), Decision Trees (DT), and Random Forests (RF) classifiers is assessed using both traditional and probabilistic performance measures. Two classification performance measures were selected: accuracy and the MCP curve (Multiclass Classification Performance) [12, 13]. The choice of the MCP curve is motivated by the fact that this measure is calculated from the probabilities returned by the classifier. Unlike $Acc$ or other popular measures such as F–score or Matthews correlation coefficient, the area under the MCP curve implicitly considers the probabilistic nature of predictions.
@@
Tables 2, 3, 4 and 5 show, for each dataset (row), several columns: Dataset name, accuracy ($Acc$), probabilistic accuracy ($Acc^{\star}$), certainty accuracy ($Acc_{v}^{\star}$), uncertainty accuracy ($Acc_{u}^{\star}$), divergence ($div$) in percentage, certainty ratio ($\mathcal{C}_{\rho}$) in percentage, and the area under the MCP curve (MCP).
@@
The 3–Nearest Neighbor classifier exhibits stable behavior, with divergence of about 4.5, resulting in a small difference between $Acc=0.809$ and $Acc^{\star}=0.783$. However, the certainty ratio $\mathcal{C}_{3-NN}=73.5\%$ due to the high value of $Acc_{u}^{\star}$, i.e., the classifier makes correct predictions by chance based on uncertainty.
@@
The Naïve–Bayes classifier shows similar behavior to the 3–NN classifier but with generally lower accuracy and uncertainty. The mean accuracy is $Acc=0.709$, and the certainty ratio is 65.8%. This suggests that Naïve–Bayes often provides confident predictions, but these predictions are not always as accurate, because the contribution of uncertainty to accuracy is high ($Acc_{u}^{\star}$=0.390).
@@
The Random Forest classifier achieves the highest accuracy ($Acc=0.856$), suggesting it is the best classifier. However, it also has a relatively high divergence of 8.6 and a certainty ratio of 81.9%, indicating that a significant portion of its performance comes from uncertain predictions. This implies that despite high accuracy, RF may not always be the most reliable.
@@
In contrast, the Decision Tree classifier exhibits a very high certainty ratio of 96.2%, because the contribution of uncertainty to accuracy is very low ($Acc_{u}^{\star}=0.058$) indicating robust and reliable performance. Although its accuracy ($Acc=0.799$) is lower than 3–NN and RF, its almost inexistent divergence and very high certainty ratio make it a more dependable classifier. The Decision Tree’s results for the area under the MCP curve are better than those of 3–NN, Naïve–Bayes and Random Forest, underscoring its reliability.
@@
The IMCP curve, which considers probabilistic predictions, consistently shows values closer to $Acc^{\star}$ than to $Acc$, supporting the argument that probabilistic measures offer a more realistic assessment of classifier performance.
@@
In summary, relying solely on accuracy is not advisable when the certainty ratio is not high (close to 100%). More broadly, we should avoid relying on any classification performance measure when the certainty ratio indicates a significant contribution of uncertainty in the confusion matrix. Therefore, it is more appropriate to analyze the probabilistic confusion matrix to assess the reliability of a performance measure, guided by the certainty ratio $\mathcal{C}_{\rho}$. This approach ensures that the evaluation of a classifier accounts not only for overall performance but also for the level of confidence and certainty in its predictions. This aspect is critical in many application realms, e.g., in clinical diagnosis.
@@
3–NN 0.809 0.783 0.829 0.304 4.9 73.5 0.720
Decision Tree 0.799 0.799 0.800 0.058 0.1 96.2 0.757
Naive Bayes 0.709 0.689 0.723 0.390 4.3 65.8 0.652
Random Forest 0.856 0.761 0.877 0.189 8.6 81.9 0.661